OpenStudy (anonymous):

Sergio's internet provider charges its customers $8 per month plus 5¢ per minute of on-line usage. Sergio received a bill from the provider covering a 2-month period and was charged a total of $46.50. How many minutes did he spend on-line during that period? (Round to the nearest whole minute, if necessary.)

OpenStudy (tkhunny):

What say you? No minutes: $8 1 minute: $8 + 0.05 = $8.05 2 minutes: $8 + 0.05*2 = $8.10 ... 10 minutes: $8 + 0.05*10 = $8.50 ... 100 minutes: $8 + 0.05*100 = $13.00 ... How far shall we go to get $46.50?

OpenStudy (anonymous):

$8 + .05*770 = $46.50

OpenStudy (anonymous):

Would it 770 minutes? @tkhunny

OpenStudy (tkhunny):

Works for me. Why do you doubt? Good work.

OpenStudy (anonymous):

Just checked my teacher's answer key and it says 610 minutes. But I don't see how it could be!

OpenStudy (tkhunny):

Check it out. 8.00 + 610*0.05 = 8.00 + 30.05 = 38.05 -- Nope. Go ahead and prove it. No need to doubt. :-)

OpenStudy (anonymous):

Thanks!